# Homogeneous potentials, Lagrange’s identity and Poisson geometry

### v2

@@ 1.1 Basic tensor invariants of Hamiltonian systems
2 Homogeneous potentials
@@ 3 Second family of potentials obtained
4 Euler-type vector fields
Let us consider homogeneous Hamilton functions $H(x)$
$H(\lambda x)=\lambda^{k}H(x)$
on the phase space $M$ with coordinates $x=(x_{1},\ldots,x_{n})$ equipped with the homogeneous Poisson bivector $P(x)$
$P(\lambda x)=\lambda^{\kappa}P(x)\,$,
where $k$ and $\kappa$ could be different We suppose that bivector
$P^{\prime}=X\wedge Y\,,\qquad X=PdH\,,\qquad Y=\sum x_{i}\frac{\partial}{\partial x_{i}}$
is an invariant rank two Poisson bivector
$\mathcal{L}_{X}P^{\prime}=0\,,\qquad[\,[P^{\prime},P^{\prime}]\,]=0\,.$
Here $Y$ is the so-called Euler vector field [16], which does not the tensor invariant of Hamiltonian flow
$\mathcal{L}_{X}Y\neq 0\,,\qquad\mathcal{L}_{X}Y\wedge X=0\,.$
The sum
$\tilde{P}=P^{\prime}-\alpha(k,\kappa)HP$
is an invariant Poisson bivector at the special value of $\alpha(k,\kappa)$, rank $\tilde{P}$=rank $P$.
$\{s_{i},s_{j}\}=\varepsilon_{ijk}s_{k}\qquad\mbox{and}\qquad\{t_{i},t_{j}\}=\varepsilon_{ijk}t_{k}\,,\qquad i,j,k=1,2,3.$ (4.5)
Here $\varepsilon_{ijk}$ is the Levi-Civita symbol.
Proposition 5
Associated with homogeneous quadratic Hamiltonian
$H=\sum_{i,j=1}^{6}a_{ij}x_{i}x_{j},\qquad x=(s_{1},s_{2},s_{3},t_{1},t_{2},t_{3})$
and homogeneous bivector $P$ (4.5) vector field $X=PdH$ together with the Euler vector field
$Y=\sum x_{i}\frac{\partial}{\partial x_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=X\wedge Y\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-HP$
has a rank four.
The proof consists of a direct verification.
$\{L_{i},L_{j}\}=\varepsilon_{ijk}L_{k},\qquad\{L_{i},p_{j}\}=\varepsilon_{ijk}p_{k},\qquad\{p_{i},p_{j}\}=0\,.$ (4.6)
For homogeneous quadratic Hamiltonian $H$ on $e^{*}(3)$ we have an analog of the previous Proposition.
Proposition 6
Associated with Hamiltonian
$H=\sum_{i,j=1}^{2}a_{ij}L_{i}L_{j}+b_{1}x_{1}+b_{2}x_{2}+b_{3}x_{3},$
and homogeneous bivector $P$ (4.6) vector field $X=PdH$ together with the Euler-type vector field
$Y=\frac{1}{2}\sum_{i=1}^{3}L_{i}\frac{\partial}{\partial L_{i}}+\sum_{i=1}^{3}p_{i}\frac{\partial}{\partial p_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=X\wedge Y\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-\frac{1}{2}HP$
has a rank four.
The proof consists of a direct verification.

### v3

@@ 1.1 Basic tensor invariants of Hamiltonian systems
1.2 Invariant Volterra brackets
In 1895-1899 Vito Volterra proposed a formal Hamiltonian description of second-order homogeneous vector fields with two quadratic integrals of motion. Our aim is to present similar formal Hamiltonian description for homogeneous vector fields with one homogeneous integral of motion.
In [15] Volterra consider dynamical system
$\dot{x}_{k}=\sum a_{k}^{ij}x_{i}x_{j}\,,\qquad a_{k}^{ij}\in\mathbb{R}\,.$ (1.12)
with two quadratic first integrals
$F_{A}=\frac{1}{2}(Ax,x),\qquad F_{B}=\frac{1}{2}(Bx,x).$
If quadratic form $F_{A}$ is positive-definite and $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different real roots of the characteristic polynomial
$\mbox{det}(B-\lambda A)=0,$
then a linear change of variables reduces the system of differential equations (1.12) to the form
$\dot{x}_{k}=\sum_{i<j\,i,j\neq k}c^{k}_{ij}(\lambda_{i}-\lambda_{j})x_{i}x_{j}$ (1.13)
where
$c^{j}_{ki}=c^{i}_{jk}=c^{k}_{ij}=-c^{k}_{ik}=-c^{i}_{kj}=-c^{i}_{jk},$
and other $c^{k}_{ij}$ are equal to zero. The Volterra proof is based on reduction of the pair of quadratic forms $F_{A}$ and $F_{B}$ to canonical form
$F_{A}=\frac{1}{2}\left(x^{2}_{1}+\cdots+x^{2}_{n}\right),\qquad F_{B}=\frac{1}{2}\left(\lambda_{1}x^{2}_{1}+\cdots+\lambda_{n}x^{2}_{n}\right).$
Then Volterra represents equations (1.13) by using the Jacobian of the first integrals
$\dot{x}_{k}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\frac{\partial(F_{A},F_{B})}{\partial(x_{i},x_{j})}$
and by using almost Poisson bivectors
$X=P_{A}dF_{B}=P_{B}dF_{A}\,.$ (1.14)
Here bivectors $P_{A}$ and $P_{B}$ is defined by the following almost Poisson brackets between coordinates
$\{x_{i},x_{j}\}_{A}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}x_{k}\,,\qquad\{x_{i},x_{j}\}_{B}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\lambda_{k}x_{k}$
which does not satisfy the Jacobi identity in the general case.
In modern terms Volterra proposed formal Hamiltonian description (1.14) for the second order homogeneous vector field (1.12). Using the Jacobi last multiplier theory he also proved that the phase flow of the formal Hamiltonian system (1.14) preserves the volume $n$-form
$dx_{1}\wedge\cdots\wedge dx_{n}$
if all the numbers $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different. Discussion of the Volterra formal Hamiltonian description (1.14) may be found in [17, 18].
Some examples of differential equations of the form (1.14) in three-dimensional space were considered by Volterra in connection with studying the dynamics of the Earth’s poles, see collection of papers [16]. Recently these Volterra calculation was repeated for other dynamical systems in three-dimensional space [19].
Let us consider smooth vector field
$V=\sum_{i=1}^{n}V_{i}(x_{1},\ldots,x_{n})\frac{\partial}{\partial x_{i}}$
with the homogeneous components
$V_{i}(\lambda x_{1},\ldots,\lambda x_{n})=\lambda^{\alpha}V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n\,.$
The Lie bracket between $V$ and the Euler vector field $U$ is equal to
$V,U]=(\alpha-1)V\,,\qquad U=\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial x_{i}}\,.$
according to the Euler theorem
$\sum_{k=1}^{n}x_{k}\frac{\partial V_{i}(x_{1},\ldots,x_{n})}{\partial x_{k}}=\alpha V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n.$ (1.15)
Bivector
$P^{\prime}=V\wedge U$ (1.16)
is a tensor invariant of the phase flow associated with $V$ since
$\mathcal{L}_{V}\,P^{\prime}=[V,V]\wedge U+V\wedge[V,U]=0\,.$
Here $\mathcal{L}_{V}\,P^{\prime}$ is a Lie derivative of the tensor field $P^{\prime}$ along the vector field $V$. Moreover, bivector $P^{\prime}$ (1.16) satisfies to the Jacobi condition since
$P^{\prime},P^{\prime}]\!]=2V\wedge[U,V]\wedge U=0\,$,
where [[.,.]][\![.,.]\!] is a Schouten-Nijenhuis bracket.
Let us suppose that function $F(x)$ is a homogeneous first integral of the vector field $V$, i.e.
$V(F)=\sum_{i=1}^{n}V_{i}\frac{\partial F}{\partial x_{i}}=0\qquad\mbox{and}\qquad U(F)=\sum_{i=1}^{n}x_{i}\frac{\partial F}{\partial x_{i}}=\kappa F\,.$
In this case we have
$P^{\prime}dF=-V(F)\,U+U(F)\,V=\kappa F\,V\,$,
that allows us to represent homogeneous vector field $V$ in the formal Hamiltonian form
$V=P^{\prime}dH\,,\qquad H=\kappa^{-1}\ln|F|\,.$ (1.17)
Here $P^{\prime}$ is the rank two Poisson bivector (1.16) and $H$ is a Hamilton function.
We suppose that if $V=PdH$ is a Hamiltonian vector field defined by homogeneous $P$ and $H$, the sum
$\tilde{P}=P^{\prime}-\alpha(k,\kappa)HP$
is an invariant Poisson bivector at the special value of $\alpha(k,\kappa)$, rank $\tilde{P}$=rank $P$, i.e.
$\tilde{P},\tilde{P}]\!]=2\alpha[\![P^{\prime},HP]\!]+\alpha^{2}[\![HP,HP]\!]=2\alpha[\![PdH\wedge U,HP]\!]+\alpha^{2}[\![HP,HP]\!]=0\,.$
$\{s_{i},s_{j}\}=\varepsilon_{ijk}s_{k}\qquad\mbox{and}\qquad\{t_{i},t_{j}\}=\varepsilon_{ijk}t_{k}\,,\qquad i,j,k=1,2,3,$ (1.18)
and the corresponding homogeneous bivector $P_{s}$. Here $\varepsilon_{ijk}$ is the Levi-Civita symbol.
Proposition 1
Associated with homogeneous quadratic Hamiltonian
$H=\sum_{i,j=1}^{6}a_{ij}x_{i}x_{j},\qquad x=(s_{1},s_{2},s_{3},t_{1},t_{2},t_{3})$
and homogeneous bivector $P_{s}$ (1.18) vector field $V=P_{s}dH$ together with the Euler vector field
$U=\sum x_{i}\frac{\partial}{\partial x_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-HP$
has a rank four.
The proof consists of a direct verification.
$\{L_{i},L_{j}\}=\varepsilon_{ijk}L_{k},\qquad\{L_{i},p_{j}\}=\varepsilon_{ijk}p_{k},\qquad\{p_{i},p_{j}\}=0\,$ (1.19)
and the corresponding homogeneous bivector $P_{e}$. For homogeneous quadratic Hamiltonian $H$ on $e^{*}(3)$ we have an analog of the previous Proposition which can be extended for the case of quasi-homogeneous Hamiltonian.
Proposition 2
Associated with quasi-homogeneous Hamiltonian
$H=\sum_{i,j=1}^{2}a_{ij}L_{i}L_{j}+b_{1}p_{1}+b_{2}p_{2}+b_{3}p_{3},$
and bivector $P_{e}$ (1.19) vector field $V=P_{e}dH$ together with the Euler-type vector field
$U^{\prime}=\frac{1}{2}\sum_{i=1}^{3}L_{i}\frac{\partial}{\partial L_{i}}+\sum_{i=1}^{3}p_{i}\frac{\partial}{\partial p_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U^{\prime}\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-\frac{1}{2}HP$
has a rank four.
The proof consists of a direct verification.
2 Homogeneous potentials on $T^{*}\mathbb{R}^{n}$